hennessy:

Solve 6 over x minus 3 equals 3 over x for x and determine if the solution is extraneous or not

lowkey:

\[\frac{ 6 }{ x }-3=\frac{ 3 }{ x }\]

lowkey:

Do you see how I got this equation?

lowkey:

Since it is the same denominator, we can solve for like terms. \[\frac{ 6 }{ x }-3=\frac{ 3 }{ x }\] \[\frac{ 6 }{ x }-\frac{ 6 }{ x }-3=\frac{ 3 }{ x }-\frac{ 6 }{ x }\]

lowkey:

Correct? Sorry for being slow, the equation was hard to type

hennessy:

the answer was x=-3, but I don't know how to do the extraneous part

lowkey:

\[3=-\frac{ 3 }{ x }\] \[3 \times x=-\frac{ 3 }{ x }\times x\] \[3x=-3\] \[x=-1\]

lowkey:

x was negative one. Do you see your mistake?

hennessy:

that's not an option

lowkey:

Wait really lol

lowkey:

My apologies

lowkey:

x would be positive one. I forgot to carry the negative.

lowkey:

Is that an option?

hennessy:

lowkey:

That makes sense. I thought the minus three was not apart of the fraction. Let me redo it.

lowkey:

 For step one, you can get rid of the denominator for the products side. Step two, getting rid of the denominator for the left hand side. Agree?

lowkey:

6x=3x-9 -3x -3x 3x=-9 x=-3

lowkey:

And to figure out if the solution is extraneous, we will plug it in the equation given. 

lowkey:

So your answer is D.

hennessy:

yeah that's what I got lol

hennessy:

okay thankyou somuch

lowkey:

You're welcome!